### 18 Mar 2026 v1

# Generative Control as Optimization: Time Unconditional Flow Matching for Adaptive and Robust Robotic Control

Abstract
Diffusion models and flow matching have become a cornerstone of robotic imitation learning, yet they suffer from a structural inefficiency where inference is often bound to a fixed integration schedule that is agnostic to state complexity. This paradigm forces the policy to expend the same computational budget on trivial motions as it does on complex tasks. We introduce Generative Control as Optimization (GeCO), a time-unconditional framework that transforms action synthesis from trajectory integration into iterative optimization. GeCO learns a stationary velocity field in the action-sequence space where expert behaviors form stable attractors. Consequently, test-time inference becomes an adaptive process that allocates computation based on convergence—exiting early for simple states while refining longer for difficult ones. Furthermore, this stationary geometry yields an intrinsic, training-free safety signal, as the field norm at the optimized action serves as a robust out-of-distribution (OOD) detector, remaining low for in-distribution states while significantly increasing for anomalies. We validate GeCO on standard simulation benchmarks and demonstrate seamless scaling to $\pi_{0}$-series Vision-Language-Action (VLA) models. As a plug-and-play replacement for standard flow-matching heads, GeCO improves success rates and efficiency with an optimization-native mechanism for safe deployment. Video and code can be found at https://hrh6666.github.io/GeCO/.
Machine Learning, ICML
1 Introduction
Generative modeling has rapidly established itself as a cornerstone of modern robotic imitation learning (Black et al., 2024; Chi et al., 2025; Janner et al., 2022; Huang et al., 2025b, a; Ze et al., 2024; Kim et al., 2024; Bu et al., 2025). By representing policies as conditional distributions rather than deterministic mappings, approaches such as diffusion models (Ho et al., 2020; Song et al., 2020; Zhang et al., 2023) and flow matching (Lipman et al., 2022; Liu et al., 2022) have demonstrated exceptional capabilities in capturing the multi-modal nature of human behavior. These methods excel at synthesizing complex trajectories for manipulation and locomotion tasks (Chi et al., 2025; Intelligence et al., 2025; Huang et al., 2025b; Ze et al., 2024; Ajay et al., 2022), driving state-of-the-art performance across diverse benchmarks (Chen et al., 2025b; Liu et al., 2023a; Zhang et al., 2025). Their success stems largely from iteratively refining noise into feasible actions via time-conditioned dynamics and effectively decomposing the challenging generation problem into a sequence of manageable denoising or flow integration steps (Ho et al., 2020; Lipman et al., 2022).
Yet, this reliance on time-conditioning introduces a fundamental structural inefficiency for control applications, a limitation that persists across standard formulations. These methods learn dynamic vector fields that evolve according to a pre-defined fictitious time schedule (Wang and Du, 2025): while this schedule bridges Gaussian noise and the expert data distribution, it creates a rigid inference paradigm that is agnostic to the actual complexity of the current robot state. The policy is compelled to execute a fixed number of integration steps regardless of convergence, thereby wasting computational resources on trivial motions while potentially failing to sufficiently refine actions for complex tasks. This “blind integration” problem not only hinders efficiency but also obscures the geometric structure of the policy. Time-varying fields lack a stationary energy landscape (Wang and Du, 2025; Sun et al., 2025): with the field direction shifting across timesteps, there is no intrinsic mechanism to verify action validity or detect out-of-distribution (OOD) scenarios, a critical gap for safe real-world deployment (Hodge et al., 2025).
To address these limitations, we reimagine generative control through a time-unconditional lens, grounding action synthesis in iterative optimization rather than trajectory integration. Our framework, Generative Control as Optimization (GeCO), embodies a paradigm shift: instead of learning time-dependent dynamics, we learn a single, stationary velocity field in the action-sequence space where expert behaviors form stable attractors. To ensure robust convergence in continuous action spaces, we incorporate a velocity rescaling mechanism that modulates flow magnitude based on distance to the expert manifold, creating a geometric sink around target modes. This transforms inference from a fixed-schedule process into an adaptive optimization loop that naturally settles at equilibrium.
Figure 1: Generative Control as Optimization (GeCO). (A) The Paradigm Shift: Unlike standard flow matching which relies on rigid, time-dependent integration schedules (top), GeCO learns a stationary velocity field where inference becomes an iterative optimization process toward stable attractors (bottom). (B) Adaptive Computation: This formulation enables the policy to dynamically allocate computational budget based on state complexity—exiting early for simple transit phases (Scenario 1) while performing deep refinement for precise manipulation (Scenario 2). (C) Intrinsic Safety: The stationary geometry provides a zero-shot safety mechanism. In-distribution (ID) states converge to low-energy equilibria ($f_{\theta}||\approx 0$), whereas out-of-distribution (OOD) anomalies exhibit persistently high field norms ($f_{\theta}||\gg 0$), enabling robust detection.
The stationary nature of GeCO’s velocity field unlocks two transformative capabilities absent in time-dependent baselines, directly addressing our core motivations of adaptive computation and intrinsic OOD awareness. First, it enables adaptive inference rooted in convergence: rather than adhering to an arbitrary step count, the policy dynamically adjusts computation by exiting early for simple states (e.g., basic reaching) and spending more time refining actions for high-precision tasks (e.g., delicate manipulation). This decouples planning horizon from training schedules, allocating computational resources efficiently. Second, and crucially, the static geometric structure provides intrinsic OOD awareness as a training-free safety signal. The residual norm of the velocity field acts as a robust epistemic uncertainty metric: in-distribution states converge to zero-velocity attractors (low norm), while OOD states, lacking a learned manifold to settle on, exhibit persistent, non-vanishing gradients. This allows the robot to detect anomalies directly from the field’s energy, without auxiliary networks or ensembles.
Notably, GeCO exhibits strong architectural versatility, seamlessly supporting both standard Diffusion Transformers (DiT) (Chi et al., 2025; Dong et al., 2024, 2025) and large-scale Vision-Language-Action (VLA) foundation models (Black et al., 2024; Intelligence et al., 2025). A defining advantage lies in its plug-and-play design: it serves as a direct drop-in replacement for standard flow-matching heads (Black et al., 2024), requiring minimal architectural adjustments or modifications to the loss function to integrate into existing systems. Building on this versatility, we conduct extensive validation of GeCO on standard simulation benchmarks (Chen et al., 2025b; Liu et al., 2023a; Zhang et al., 2025). Empirically, GeCO outperforms state-of-the-art time-conditioned baselines while fundamentally redefining inference efficiency through the modulation of computational resources based on task complexity, rather than adherence to rigid fixed schedules. Beyond pure performance gains, we further establish that the intrinsic geometric signals of GeCO’s stationary field enable robust detection of task-level distribution shifts, delivering an optimization-native safety mechanism critical for reliable real-world robotic deployment.
2 Related Work
Generative Models for Robotic Control. Generative policy learning has become a dominant paradigm for robotic control in recent years. One prominent direction leverages the reasoning and perceptual capabilities of Vision-Language Models (VLMs) (Beyer et al., 2024; Bai et al., 2023; Liu et al., 2023b) for high-level planning (Driess et al., 2023; Liang et al., 2022; Tian et al., 2024) and action generation (Kim et al., 2024, 2025; Bu et al., 2025). Parallel efforts utilize video generation models (Liu et al., 2024b; Yang et al., 2024; Wiedemer et al., 2025) to serve as predictive priors for robotic decision-making (Du et al., 2023; Chen et al., 2025a; Shen et al., 2025). In the realm of precise robotic control, diffusion (Ho et al., 2020; Song et al., 2020) and flow matching (Lipman et al., 2022) have demonstrated exceptional effectiveness, functioning as robust planners and action-chunk generators (Chi et al., 2025; Huang et al., 2025a; Zhou et al., 2024; Huang et al., 2025b; Janner et al., 2022; Ajay et al., 2022; Dong et al., 2025; Ze et al., 2024). Notably, the integration of continuous flow-matching heads into VLA architectures has emerged as a powerful paradigm (Black et al., 2024; Intelligence et al., 2025; Li et al., 2025; Shukor et al., 2025; Jiang et al., 2025); this approach bridges the semantic understanding of VLM backbones with high-frequency, continuous action execution. Such continuous generation mechanisms are widely adopted in modern Vision-Language-Action systems to replace purely autoregressive, discretized action decoding (Kim et al., 2024, 2025), enabling higher fidelity control and improved real-time performance (Black et al., 2025; Intelligence et al., 2025). Our work focuses on this continuous generative control family of flow matching, which continues to drive advancements across domains ranging from robotic manipulation to navigation.
Time-Unconditional Generative Dynamics. Most diffusion and flow-based models (Ho et al., 2020; Song et al., 2020; Lipman et al., 2022) formulate generation as a time-conditioned dynamical system, where the vector field varies with an explicit noise or time variable, typically necessitating a fixed integration horizon. Recent studies (Sun et al., 2025) indicate that removing this explicit conditioning can lead to graceful degradation, or even performance gains, suggesting that dynamics can often be inferred implicitly from the state. However, naively discarding time conditioning does not guarantee stable equilibrium behavior or well-posed optimization dynamics (Wang and Du, 2025). (Wang and Du, 2025) address this by learning a stationary vector field compatible with an implicit energy landscape (LeCun et al., 2006), enabling optimization-driven sampling (Florence et al., 2022). While this paradigm shows promise in image generation, its effectiveness in robotic control, which demands rigorous accuracy and sequential stability, still remains an open question.
3 Method
Our goal is to turn generative control for robotic manipulation from a fixed-time integration process into an optimization problem over a time-invariant velocity field. At each control step, the policy observes images, proprioception, and a language instruction, and synthesizes a short-horizon action sequence. We first revisit time-conditioned diffusion and flow matching for such control, then introduce our time-unconditional field and optimization-based inference, and finally show how the same geometry yields intrinsic OOD awareness.
3.1 Revisiting Time-Conditioned Generative Control
Diffusion- and flow-based controllers for manipulation typically learn a time-conditioned velocity field
$v_{\theta}(x_{\gamma},\gamma,s_{t}),$
where $x_{\gamma}$ is a noisy or partially denoised action sequence, $\gamma$ is a time variable, and $s_{t}$ is the current observation and instruction. Inference follows an integration path in the time domain: starting from noise, the policy traverses $\gamma$ from 0 to $1$ using steps $\{\Delta\gamma_{k}\}_{k}$ with $\sum_{k}\Delta\gamma_{k}\approx 1.$
Although one can change the number of steps and step sizes at deployment, these choices are constrained by the requirement of representing a valid integration of the time-conditioned field over $\gamma\in[0,1]$. Stopping too early under-integrates the ODE/SDE; letting $\sum_{k}\Delta\gamma_{k}>1$ pushes the system beyond the interval on which the field is defined and trained. Thus the geometry of $v_{\theta}(\cdot,\gamma,s_{t})$ is always tied to a specific notion of time progress.
This design couples three roles into a single scalar $\gamma$: tracking progress along the generative trajectory, controlling the strength of the field, and parameterizing the integration domain. As a result, the field geometry changes with $\gamma$, and there is no single, time-invariant notion of “good actions” or a creteria whose magnitude directly measures how close a final action is to the in-distribution manifold.
GeCO decouples this dependence on time: instead of conditioning on $\gamma$, we learn a stationary field over action sequences and use it as the objective of an unconstrained optimization procedure.
3.2 Time-Unconditional Velocity Fields for Control
At timestep $t$, the robot observes $s_{t}$ (multi-view RGB of the workspace, proprioception, and a language instruction) and predicts a sequence of future low-level actions.
$a=a_{t:t+T_{a}-1}\in\mathbb{R}^{T_{a}\times d_{a}},$
Rather than a time-indexed $v_{\theta}(\cdot,\gamma,s_{t})$, GeCO learns a time-unconditional velocity field $f_{\theta}(x,s_{t})$ over action sequences. We want $f_{\theta}(\cdot,s_{t})$ to point from noisy or suboptimal sequences toward equilibrium actions that solve the current task, and to vanish near ground truth in-distribution actions, in order stop automatically during inference without an explicit time condition.
To train such a field, we introduce $\gamma$ only as a implicit training variable. Given a demonstration $a$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,I)$, we form
$x_{\gamma}=\gamma a+(1-\gamma)\varepsilon,\qquad\gamma\sim\mathcal{U}(0,1),$ (1)
and crucially do not provide $\gamma$ to the model. At each $x_{\gamma}$ we define a restoring direction
$g^{\star}(a,\varepsilon,\gamma)=(\varepsilon-a)\,c(\gamma),$ (2)
where $c(\gamma)$ is a scalar schedule that decays to zero as $\gamma\to 1$. Because $c(1)=0$, the target field vanishes at $x_{\gamma}=a$, transforming ground-truth action sequences into natural stationary equilibrium points of the learned field. Given such rescaling mechanism, the denoising process is able to automatically stop in the ground truth stationary points, effectively solving the non-equilibrium nature of the vanilla time-unconditional flow matching.
Conditioned on $s_{t}$, the loss is
$\mathcal{L}(\theta)=\mathbb{E}_{(s_{t},a)}\mathbb{E}_{\varepsilon,\gamma}\Bigl\|f_{\theta}(x_{\gamma},s_{t})-g^{\star}(a,\varepsilon,\gamma)\Bigr\|^{2}.$ (3)
Here $\gamma$ only controls the interpolation and the scale via $c(\gamma)$; since it never enters the model, $f_{\theta}(\cdot,s_{t})$ must organize all training pairs into a single consistent field for each scene. This encourages $f_{\theta}$ to approximate the gradient of an implicit potential field over action sequences whose minima lie near successful demonstrations.
3.3 Optimization-Based Inference
Once $f_{\theta}$ has been learned, inference is no longer framed as integrating over a time axis. At deployment, the controller runs in closed loop: at each timestep $t$, it observes $s_{t}$ and solves a small optimization problem in action space using $f_{\theta}(\cdot,s_{t})$.
We initialize $a^{(0)}$ from a prior $p_{\text{prior}}(\cdot\mid s_{t})$, e.g., isotropic Gaussian noise, and then perform $K$ gradient-based updates:
$a^{(k+1)}=a^{(k)}-\eta_{k}\,f_{\theta}(a^{(k)},s_{t}),$ (4)
where $\eta_{k}$ may depend on the sample or iteration. Because there is no time index, and there is no requirement that the effective “total step length” $\sum_{k}\eta_{k}$ equal a particular value. Taking more or fewer steps, or even $\sum_{k}\eta_{k}>1$, remains a natural optimization process in the same stationary field, rather than violating a time-integration constraint and leading to non-equilibrium divergence.
Convergence is monitored directly through the field:
$\|f_{\theta}(a^{(k)},s_{t})\|\leq\tau_{\text{opt}}\quad\text{or}\quad k\geq K_{\max}.$ (5)
Simple scenes (e.g., unobstructed reaches) reach a low-norm region in few iterations; cluttered or contact-rich scenes require more. GeCO behaves as an optimization-based, receding-horizon manipulation controller grounded in a learned generative field. Detailed inference pseudocode can be found in Algo.1.
Algorithm 1 Adaptive Inference (GeCO)
1:Input: Observation $s$, Max steps $K$, Tolerance $\tau$, Rate $\eta$
2:Output: Action $\hat{a}$, OOD Score $S$
[…]
4:for $k=0,\dots,K-1$ do
5:  $v_{k}\leftarrow f_{\theta}(a^{(k)},s)$
6:  if $\|v_{k}\|_{2}<\tau$ then
7:   break $\triangleright$ Adaptive early exit
8:  end if
9:  $a^{(k+1)}\leftarrow a^{(k)}-\eta\cdot v_{k}$
10:end for
[…]
12:return $\hat{a}$, $\|f_{\theta}(\hat{a},s)\|_{2}$
3.4 Intrinsic OOD Detection from Field Geometry
GeCO also provides optimization-native OOD awareness. Here OOD is defined at the level of the conditional field: training exposes the model to $(s_{t},a)$ pairs from $p_{\text{data}}(s_{t},a)$, and thus to conditional fields $f_{\theta}(\cdot,s_{t})$ induced by $s_{t}\sim p_{\text{data}}(s_{t})$. At test time, an observation $s_{t}$ is OOD if it induces a field whose geometry is inconsistent with this family; optimization on such a field need not reach a low-norm equilibrium, and this breakdown in convergence becomes an OOD signal.
Given a test observation $s_{t}$, we run optimization as in (4)–(5), obtaining an approximate equilibrium $\hat{a}(s_{t})$, and define
$\mathrm{Score}(s_{t})=\bigl\|f_{\theta}(\hat{a}(s_{t}),\,s_{t})\bigr\|.$ (6)
For in-distribution robot control states $s_{t}\sim p_{\text{data}}$, the induced field resembles those seen during training: optimization can move $x_{\gamma}$ into regions that behave like high-$\gamma$ samples, where $c(\gamma)$ is small and the field is trained to nearly vanish, so $\mathrm{Score}(s_{t})$ is low. When $s_{t}$ is OOD, the induced field is itself out-of-distribution: optimization operates in regions never trained to lead to small norm, the gradient norm at the end of the optimization process at $\hat{a}(s_{t})$ remains large. We obtain a binary OOD decision via
$\mathrm{isOOD}(s_{t})=\mathbf{1}\{\mathrm{Score}(s_{t})>\tau_{\text{OOD}}\}.$ (7)
This incurs essentially no overhead, as the same gradient norms are already computed for convergence checking.
Figure 2: Computation Follows Task Complexity. We visualize the spatial distribution of inference effort along a single rollout. The first three panels (a–c) are sampled from LIBERO-Spatial, and the last three panels (d–f) are from LIBERO-Object. The color of each line encodes the number of function evaluations (NFE) required for convergence at that state, ranging from blue (NFE = 1) to red (NFE = 20). This visualization illustrates how GeCO allocates more computation to challenging states while using fewer steps in easier regions.
3.5 Plug-and-Play Integration with VLA Systems
Modern flow-based Vision-Language-Action (VLA) systems (Black et al., 2024; Intelligence et al., 2025; Jiang et al., 2025; Shukor et al., 2025) consist of two core components: (1) a Vision-Language Model (VLM) that fuses multi-modal inputs (multi-view RGB $\mathbf{v}\in\mathbb{R}^{H\times W\times 3\times N_{v}}$, proprioception $\mathbf{p}\in\mathbb{R}^{d_{p}}$, language instruction $l\in\mathcal{L}$) into a task-aware conditional signal $s_{t}=\text{VLM}(\mathbf{v},\mathbf{p},l)$; (2) a time-conditioned flow-matching head that takes $s_{t}$ and action sequence $x_{\gamma}$ as inputs to predict $v_{\theta}(x_{\gamma},\gamma,s_{t})$.
GeCO enables plug-and-play integration with such VLA architectures via minimal modifications: we directly replace the time-conditioned flow-matching head $v_{\theta}(x_{\gamma},\gamma,s_{t})$ with our time-unconditional velocity field $f_{\theta}(x,s_{t})$ (Section 3.2), while retaining the VLM’s multi-modal fusion pipeline and the VLA system’s original input-output interface. No architectural changes to the VLM or control pipeline are required—at inference, the VLA system generates action sequences $\mathbf{a}\in\mathbb{R}^{T_{a}\times d_{a}}$ via GeCO’s adaptive optimization loop (Algorithm 1) using the same $s_{t}$ signal, leveraging the stationary geometry of $f_{\theta}$ to eliminate the rigid time-schedule constraints of baseline flow-matching heads.
4 Experiments
We evaluate GeCO along three complementary axes: (1) Adaptive Efficiency, analyzing how the optimization-based inference dynamically allocates computation based on task complexity; (2) Scalability, assessing the method’s effectiveness as a plug-and-play head for large Vision-Language-Action (VLA) models ($\pi_{0}$ series); and (3) Intrinsic Safety, verifying the reliability of the stationary field norm as a zero-shot Out-of-Distribution (OOD) detector. Implementation details can be found in Appendix A.
4.1 Policy Performance and Adaptive Computation
The Efficiency-Performance Trade-off.
Table 1 demonstrates that GeCO achieves a superior trade-off between inference speed and task success compared to fixed-schedule baselines on the LIBERO (Liu et al., 2023a) benchmark. Remarkably, GeCO with a budget of only 5 steps already outperforms Rectified Flow (20 steps), achieving a higher success rate (91.9% vs. 90.0%) while requiring 75% less compute. When the budget is increased to 20 steps, GeCO reaches a peak success rate of 93.5%. Crucially, the average NFE required to achieve this peak is only 11.6, significantly lower than the maximum budget. This indicates that GeCO’s optimizer effectively allocates computational resources, autonomously exiting early for simpler states while reserving the budget for complex ones.
Temporal Analysis: Computation Follows Complexity.
To verify that the observed variation in NFE is structurally driven by task complexity rather than randomness, we analyze the inference profile shown in Figure 2. We observe distinct optimization behaviors across different task types:
  • Motion Planning Complexity (LIBERO-Spatial): For tasks requiring spatial reorientation (Fig. 2 a-c), higher NFE concentrates heavily around bottleneck phases, such as pre-grasp alignment and object placement. In contrast, free-space transit phases converge rapidly, demonstrating that the learned vector field provides a smooth gradient in unobstructed regions.
  • Semantic Grounding (LIBERO-Object): In tasks where the robot must identify a specific target among multiple distractors based on language instructions (e.g. pick the ketchup and place it in the basket), we observe a distinct NFE spike at the episode onset (Fig. 2 d-f). This suggests that the optimizer requires more iterations to resolve the visual-semantic correspondence, effectively deliberating to lock onto the specific object defined by the instruction before committing to a trajectory.
These results confirm that GeCO acts as an adaptive controller, dynamically modulating its inference horizon based on both kinematic constraints and perceptual uncertainty.
Table 1: Performance vs. Efficiency on LIBERO. GeCO achieves higher success rates by allowing adaptive refinement. By increasing the max step budget, GeCO refines actions for complex states, yet the average compute remains low compared to fixed-schedule baselines.
Method Max Success Rate (%) Efficiency
Steps Goal Spat. Obj. Long Avg. NFE
Diff. Policy 100 (Fixed) 82.9 91.4 88.9 82.7 86.5 100.0
Rectified flow 20 (Fixed) 92.4 94.6 97.0 76.0 90.0 20.0
GeCO (Ours) 5 91.6 95.4 98.2 82.4 91.9 5.0
GeCO (Ours) 10 93.0 95.8 99.0 81.8 92.4 8.7
GeCO (Ours) 20 95.2 95.8 99.0 83.8 93.5 11.6
GeCO (Ours) 30 93.6 95.2 98.8 84.8 93.1 12.8
Table 2: Success rate (%) of different VLA methods on the RoboTwin 2.0 benchmark. Each task is evaluated under two difficulty levels: Easy and Hard. For every method we report the success rate for each task and difficulty level, together with the average performance across all five tasks. Higher values indicate better task completion ability. The best result in each column is highlighted in bold.
Method Adjust Bottle Beat Block Hammer Blocks Ranking (Size) Click Alarmclock Click Bell Avg.
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
RDT (Liu et al., 2024a) 81.0 75.0 77.0 37.0 0.0 0.0 61.0 12.0 80.0 9.0 60.0 27.0
ACT (Zhao et al., 2023) 97.0 23.0 56.0 3.0 0.0 0.0 32.0 4.0 58.0 3.0 49.0 7.0
$\pi_{0}$ (Black et al., 2024) 90.0 56.0 43.0 21.0 7.0 1.0 63.0 11.0 44.0 3.0 49.0 18.0
$\pi_{0}$ with GeCO 86.0 66.0 52.0 36.0 0.0 0.0 72.0 16.0 42.0 21.0 50.0 28.0
Table 3: VLA success rates (%) on VLABench tracks. We report track-wise success and the average over the listed tracks; higher is better.
Method Track 1 Track 2 Track 3 Track 4 Track 6 Avg.
$\pi_{0}$ 47.0 21.2 29.1 17.3 32.2 29.4
$\pi_{0.5}$ 40.6 22.6 18.0 16.1 25.6 24.5
$\pi_{0}$ with GeCO 61.0 22.0 34.0 20.0 45.0 36.0
4.2 Scalability to Vision-Language-Action Models
We evaluate the scalability of GeCO by integrating it into flow-matching-based VLA models, with the $\pi_{0}$-series models selected as a representative example. A critical advantage of GeCO is its architectural compatibility with existing time-conditional backbones. Since $\pi_{0}$ is originally trained as a time-dependent flow matching model expecting a time embedding $t\in[0,1]$, we adapt it to our time-unconditional framework by simply fixing the input time $t=0$ during fine-tuning. This effectively freezes the time channel, forcing the network to learn a single, stationary velocity field $f_{\theta}(x,s)$ using the pre-trained weights, rather than a sequence of time-varying fields. This plug-and-play adaptation integrate the rich visual-semantic representations of foundation models with the native adaptive optimization inference of GeCO, with no need for structural modifications or training from scratch. For a fair comparison, both GeCO and the baselines are fine-tuned from the same pre-trained $\pi_{0}$ base for 30,000 steps. Details can be found in Appendix B.
VLABench Result
We benchmark our method on VLABench (Zhang et al., 2025), a challenging suite designed to evaluate long-horizon, language-conditioned robotic manipulation with strong requirements on semantic grounding and multi-step reasoning. As shown in Table 3, $\pi_{0}$ with GeCO achieves the strongest overall performance with an average success rate of 0.36, improving upon the baseline $\pi_{0}$ (0.294). The most substantial gains are observed on Track 1 and Track 6, which respectively evaluate in-distribution manipulation skills and robustness to unseen textures and visual variations. In addition, GeCO consistently improves results on Track 3 (common-sense and world knowledge) and Track 4 (semantic instruction following), both of which emphasize multi-step reasoning and long-horizon task execution. By contrast, performance on Track 2 (cross-category generalization) remains comparable to standard $\pi_{0}$, suggesting that improvements are more pronounced on execution- and reasoning-intensive tasks than on pure category-level transfer.
RoboTwin 2.0 Benchmark Result
[…]
LIBERO Benchmark Result
LIBERO (Liu et al., 2023a) comprises four suites that isolate different distribution shifts in language-conditioned manipulation: goal changes (Goal), spatial relation shifts (Spat.), object category shifts (Obj.), and long-horizon compositional execution (Long). In Table 4, GeCO remains comparable overall relative to standard fine-tuning: for $\pi_{0}$ it matches a similar average (93.9 vs. 94.2), and for $\pi_{0.5}$ it stays close (95.9 vs. 96.9). These results suggest that GeCO maintains performance comparable to the flow matching baseline across different VLA model backbones.
Collectively, these results validate that GeCO acts as a better drop-in replacement for VLA action heads, delivering higher performance and robustness under identical training conditions.
Table 4: VLA success rates (%) on LIBERO benchmark. We report track-wise success and the average over the listed tracks; higher is better.
Method Goal Spat. Obj. Long Avg.
$\pi_{0}$ + FAST (Pertsch et al., 2025) 88.6 96.4 96.8 60.2 85.5
$\pi_{0}$ (Black et al., 2024) 95.8 96.8 98.8 85.2 94.2
$\pi_{0.5}$ (Intelligence et al., 2025) 98.0 98.8 98.2 92.4 96.9
Ours ($\pi_{0}$) 96.4 96.8 97.0 85.4 93.9
Ours ($\pi_{0.5}$) 96.6 97.0 98.8 91.2 95.9
4.3 Intrinsic OOD Detection from Optimization
GeCO performs action synthesis via iterative optimization at each planning call. We hypothesize that under distribution shift, this optimization fails to converge cleanly, leading to larger update magnitudes. We therefore use the final update norm produced by a single planning call as an intrinsic anomaly score.
Experiment Setup.
We train a single policy on LIBERO-Goal (10 tasks) and evaluate OOD detection on LIBERO-Spatial (10 tasks) as Task OOD. We choose LIBERO-Goal for training because tasks share the same scene layout and objects but differ in goal predicates, making performance sensitive to correctly grounding language instructions into behaviors. At test time, LIBERO-Spatial introduces unseen spatial arrangements of otherwise similar objects, creating a controlled workspace-layout shift while keeping the manipulation vocabulary comparable.
Baseline OOD Signal.
Following the Diff-DAgger-style idea of using the diffusion training objective as an uncertainty signal (Lee et al., 2025), we define a lightweight baseline based on the flow-matching training loss. The original formulation estimates an expectation over many sampled timesteps (e.g., 512 samples); to avoid extra sampling cost, we use a single-step proxy and evaluate the flow-matching loss at a fixed timestep $t=1$ for each planning call. This produces one scalar baseline score per planning call.
Anomaly Score and AUROC Computation.
An episode consists of multiple planning calls (each call runs optimization to produce an action chunk that is then executed in the environment). We compute an anomaly score per planning call. For the $j$-th planning call at state $s^{(j)}$, we run GeCO optimization for at most $K_{\max}=10$ steps. With the update $a_{k+1}\leftarrow a_{k}-\eta\,f_{\theta}(a_{k},s^{(j)})$, we define
$\text{score}\bigl(\text{plan }j\bigr)\;=\;\bigl\|f_{\theta}(a^{(j)}_{K},s^{(j)})\bigr\|_{2},$
where $K\leq K_{\max}$ denotes the last executed refinement step (if early termination is triggered, $K$ is the stopping step; otherwise $K=K_{\max}$). We sample 500 in-distribution (ID) episodes from LIBERO-Goal and 500 OOD episodes from LIBERO-Spatial. To avoid overweighting longer episodes, we compute AUROC on a fixed number of planning calls per episode, treating OOD plans as positives and ID plans as negatives.
Temporal Filtering and Early Reporting.
Within each planning call, raw per-step norms can be noisy, so we evaluate two lightweight filters over the refinement-step residual norms $r_{k}=\|f_{\theta}(a_{k},s^{(j)})\|_{2}$. Both filters output an online OOD flag for the current planning call; once the flag triggers (at any planning call), we immediately terminate the episode to avoid executing the remaining (eventually failing) interactions.
  • Moving Average. We smooth $r_{k}$ by averaging over the last $w=5$ refinement steps, $\tilde{r}_{k}=\frac{1}{w}\sum_{i=k-w+1}^{k}r_{i},$ and trigger an OOD flag when $\tilde{r}_{k}$ exceeds a threshold.
  • Leaky Bucket. We maintain an accumulator that integrates only sustained large residuals: $b_{k}\leftarrow\max\{0,\;(1-\lambda)b_{k-1}+(r_{k}-\tau)_{+}\},$ where $(x)_{+}=\max(x,0)$. We raise an OOD flag if $b_{k}\geq B$ at any step.
Quantitative Results.
Table 5 reports separability and operating-point performance for distinguishing ID (LIBERO-Goal) from OOD (LIBERO-Spatial). We treat OOD plans as positives and ID plans as negatives; thus TPR measures the fraction of OOD instances correctly flagged, while TNR measures the fraction of ID instances correctly retained (i.e., not falsely flagged). The baseline loss signal is less aligned with the suite-level shift, whereas GeCO yields strong separation. We choose the detection threshold to target TPR $\approx 90\%$ for our detector when possible, and report the corresponding TNR.
Time Saved measures the fraction of interaction time avoided on OOD episodes by early reporting: for each OOD episode, we record the first reporting time $t_{\mathrm{report}}$ (the environment time step when the first planning call triggers the OOD flag) and normalize by the full episode duration $T_{\mathrm{total}}$ under the standard evaluation protocol, $\text{Time Saved}=1-t_{\mathrm{report}}/T_{\mathrm{total}}$; we then average this quantity over OOD episodes.
Table 5: OOD Detection on LIBERO Suite Shift (Goal $\rightarrow$ Spatial). We report planning-level AUROC for separating ID (Goal) and OOD (Spatial) using optimization-dynamics signals computed per planning call. At a fixed operating point, TNR is specificity on ID instances and TPR is recall on OOD instances. Time Saved is the average fraction of interaction time avoided on OOD episodes by early reporting, computed as $1-t_{\mathrm{report}}/T_{\mathrm{total}}$.
Method AUROC ($\uparrow$) TNR(%) TPR (%)
Time Saved(%)
Baseline (Moving Avg) 0.53 14.4 27.8 10.3
Ours (Moving Avg) 0.93 82.4 89.7 42.4
Ours (Leaky Bucket) 0.93 84.0 90.0 44.3
4.4 Ablation Study: Effect of Truncation Parameter $\alpha$
To evaluate the sensitivity of our adaptive optimization dynamics to the truncation schedule, we conduct an ablation study on the truncation parameter $\alpha$. As discussed in Section 3.2, $\alpha$ controls the onset of the velocity rescaling decay, directly impacting the equilibrium convergence. Table 6 presents the success rates on the LIBERO benchmark across different values of $\alpha\in\{0.6,0.8,0.9\}$.
The results highlight a clear trade-off governed by the smoothness of the learned vector field. When $\alpha$ is too small (e.g., $\alpha=0.6$), the field becomes overly smooth. While this provides stable gradient directions, it drastically slows down convergence. The optimization often fails to reach the tight equilibrium manifold within a limited step budget. Conversely, an excessively large $\alpha$ (e.g., $\alpha=0.9$) leads to a large Lipschitz constant ($L$-smoothness) for the velocity field. This creates a sharp and stiff optimization landscape, making it difficult for the iterative solver to converge stably, thus degrading overall performance. Setting $\alpha=0.8$ strikes the optimal balance, ensuring the attractors are sharp enough for precise control while maintaining a smooth enough field for rapid and stable inference.
Table 6: Ablation Study on Truncation Parameter $\alpha$. We report the success rates (%) on the LIBERO benchmark across different values of $\alpha$ to demonstrate its impact on policy performance.
$\alpha$ Success Rate (%)
Goal Spat. Obj. Long Avg.
$0.6$ 85.8 98.0 91.6 48.6 81.0
$0.8$ 96.4 97.4 98.6 92.4 96.2
$0.9$ 95.8 97.8 94.0 81.2 92.2
4.5 Real-World Experiment
To validate the practical effectiveness and efficiency of GeCO in unstructured physical environments, we conduct real-world benchmarks using the Galaxea R1 Lite(Jiang et al., 2025) mobile manipulator.
Figure 3: GeCO policy execution for the Nut Assembly task. The robot performs high-precision alignment and rotational insertion.
Figure 4: GeCO policy execution for the Chemistry Tube Arrangement task. The policy adaptively handles the tight-tolerance insertion of fragile tubes.
Experimental Setup.
As shown in Fig.3 and Fig.4, we use Galaxea G0 Plus (Jiang et al., 2025) as our baseline, which employs standard flow matching (FM) as VLA action head, and compare it against G0 Plus with GeCO. For both policies, the flow-matching heads are fine-tuned on a demonstration dataset collected via teleoperation. We select two complex and high-precision manipulation tasks for the benchmark: Nut Assembly and Chemistry Tube Arrangement. These tasks require millimeter-level precision, providing a rigorous testbed for GeCO’s ability to perform deep refinement during contact-rich bottleneck phases while exiting early during simple transit motions. Each task is evaluated over 10 independent trials on the physical hardware. We control the same total gradient steps during training. Detailed experimental setup can be found in Appendix E.
Quantitative Results.
As shown in Table 7, GeCO consistently outperforms the standard flow-matching baseline in success rate across both challenging tasks, and notably achieves these gains with a significantly lower average Number of Function Evaluations (NFE). While the baseline is constrained to a fixed schedule of 10 steps, GeCO’s adaptive inference mechanism allows it to settle at equilibrium faster, thereby improving efficiency without compromising control fidelity (see Appendix F for detailed NFE analysis). We attribute this substantial performance margin to two key advantages: enhanced sample efficiency and superior generalization. First, given that both policies were fine-tuned using identical datasets and training budgets, GeCO’s higher success rate clearly underscores the improved sample efficiency inherent to its time-unconditional formulation. Second, our method demonstrates remarkable robustness during physical deployment. Under domain-randomized evaluations, GeCO’s optimization-based inference maintains robust convergence amidst real-world visual and physical variations, whereas the fixed-schedule baseline proves brittle to such unstructured domain shifts.
Table 7: Real-world performance on Galaxea R1 Lite. We report the success rate and the average NFE over 10 trials per task.
Method Nut Assembly Tube Arrangement
Success ($\uparrow$) Avg. NFE Success ($\uparrow$) Avg. NFE
G0 Plus (Baseline) 10% 10.0 (Fixed) 20% 10.0 (Fixed)
G0 Plus + GeCO (Ours) 70% 6.2 80% 5.1
5 Discussion and Limitations
GeCO reformulates generative action synthesis as a robust optimization task, and the proposed framework delivers promising performance alongside strong adaptivity and OOD awareness, which validates the effectiveness of our core design. While the framework performs well in practical settings, we acknowledge that its theoretical foundation can be further strengthened. Future work will focus on deriving more rigorous theoretical bounds for performance gains, the convergence speed of the optimization process, and the OOD detection mechanism in the context of control and VLA to further solidify GeCO’s theoretical foundation.
6 Conclusion
We introduced GeCO, a time-unconditional formulation of flow-matching policies that casts action generation as iterative optimization over a stationary velocity field, where expert behaviors form stable attractors via velocity rescaling. This removes the rigid dependence on fixed integration schedules and provides a plug-and-play replacement for continuous action heads in modern VLA systems. Experimentally, GeCO delivers a better efficiency–performance trade-off through state-dependent adaptive optimization, scales to flow matching based VLA models, and yields an optimization-native safety signal: the final field norm serves as a training-free OOD score.
Impact Statement
This paper aims to advance the efficiency and safety of generative robotic control. The proposed framework, GeCO, introduces an intrinsic anomaly detection mechanism that helps mitigate operational risks in unstructured environments. Additionally, our adaptive inference approach optimizes computational resource usage, enabling more efficient deployment on robot hardware. We do not foresee specific negative societal consequences beyond the general implications of advancing automation technology.
Appendix A Implementation Details for Adaptive Inference and Training
A.1 Codebase and Baselines
All experiments are implemented on top of the CleanDiffuser (Dong et al., 2024) codebase. We use the same network backbone as a Continuous Rectified Flow policy (Liu et al., 2022), but differ in the learning objective and the test-time inference algorithm.
A.2 Model Architecture
We follow the Rectified Flow architecture with a 1D diffusion-style transformer and a frozen vision-language conditioner.
Vision-language condition.
We use a frozen DINOv2 (Oquab et al., 2023) vision encoder and a T5 (Raffel et al., 2020) text encoder. The T5 hidden dimension is set by the pretrained configuration:
$\texttt{t5\_hidden\_dim}\leftarrow\texttt{T5Config.from\_pretrained(t5\_model).d\_model}.$
The conditioner is instantiated as:
  • ViTAndT5VisionLanguageCondition with emb_dim=768, freeze=True, To=1, and n_views=2.
Policy backbone.
The action generator uses a DiT-style 1D transformer with cross-attention conditioning:
  • DiT1dWithACICrossAttention with x_dim=act_dim, x_seq_len=Ta(=16), emb_dim=768, d_model=384, n_heads=6, depth=12.
GeCO uses the same backbone modules for fair comparison, while changing the training objective and inference procedure.
A.3 Adaptive Inference Setup
Sampler.
We use a standard gradient-descent (GD) sampler with a maximum of $K_{\max}=30$ refinement steps. Denoting the refinement direction by $f_{\theta}(a_{k},s)$, the update is:
$a_{k+1}\leftarrow a_{k}-\eta_{k}\,f_{\theta}(a_{k},s),\quad k=0,\dots,K_{\max}-1.$
Step-size schedule.
We use a fixed per-step learning-rate schedule $\{\eta_{k}\}_{k=1}^{30}$:
$\eta_{1}=0.1,\quad\eta_{2..4}=0.05,\quad\eta_{5..16}=0.02,\quad\eta_{17..30}=0.01.$
We enable adaptive early stopping using a norm-based stopper: we terminate refinement once
$\|f_{\theta}(a_{k},s)\|_{2}<0.4.$
A.4 Training Recipe and Hyperparameters
Pretrained encoders.
  • T5: google-t5/t5-base
  • ViT: facebook/dinov2-base
Training details.
We use a truncated linear decay for equilibrium field learning
$c(\gamma)=\begin{cases}\lambda,&\gamma\leq a_{0},\\ \lambda\dfrac{1-\gamma}{1-a_{0}},&\gamma>a_{0},\end{cases}$ (8)
with scale $\lambda=4$ and onset $a_{0}=0.1$.
Constants.
  • Observation horizon: $T_{o}=1$
  • Action horizon: $T_{a}=16$
  • Executed actions per step: num_act_exec=8
  • Normalization parameters: norm_params=[0.5, 0.5, 0.5]
Appendix B Benchmarks
LIBERO.
We evaluate on the LIBERO simulation benchmark (Liu et al., 2023a), which contains four task suites—Goal, Spatial, Object, and Long—designed to isolate different distribution shifts in language-conditioned manipulation. Goal varies task goals while keeping objects and layouts fixed; Spatial introduces unseen spatial configurations; Object shifts object categories; and Long consists of long-horizon, compositional tasks. For policy learning and evaluation, our model predicts action chunks of length $T_{a}=16$; after each prediction, we execute the first $8$ actions (num_act_exec=8) in closed loop before replanning the next chunk. The observation space consists of two-view RGB images at the current time step with $T_{o}=1$ (no history), and the language instruction is encoded by a frozen T5 encoder (google-t5/t5-base). Unless otherwise specified, we train each suite-specific policy for 30k optimization steps and report success rates averaged over 50 rollouts, following the standard evaluation protocol of LIBERO.
RoboTwin 2.0.
We additionally benchmark on RoboTwin 2.0 (Chen et al., 2025b), a bimanual manipulation benchmark that evaluates robustness under controlled domain shifts. RoboTwin 2.0 reports performance under two difficulty settings: Easy uses clean environments, while Hard introduces domain-randomized perturbations such as clutter distractors, texture changes, lighting variations, and tabletop-height shifts. We evaluate on five representative tasks (Adjust Bottle, Beat Block Hammer, Blocks Ranking (Size), Click Alarmclock, and Click Bell) and report Easy/Hard success and their averages (Table 2). All methods are evaluated with 100 rollouts per setting, using the official success criteria of RoboTwin 2.0. We highlight that the Hard setting creates a realistic robustness test by perturbing both visual appearance and scene dynamics while keeping task semantics unchanged.
VLABench.
We further evaluate on VLABench (Zhang et al., 2025), a large-scale benchmark for language-conditioned manipulation that emphasizes long-horizon reasoning and semantic grounding. VLABench organizes tasks into multiple tracks covering in-distribution execution, cross-category generalization, common-sense/world-knowledge transfer, semantic instruction understanding, and unseen texture generalization. We report track-wise success rates on Tracks and their average (Table 3), following the benchmark’s standard evaluation protocol with 50 rollouts. We include per-task breakdown for the 10-task primitive subset in Table 8, which complements the track-level summary in Table 3.
Table 8: Per-task success rates on VLABench primitive 10-task evaluation. We report success rates for the baseline $\pi_{0}$ and $\pi_{0}$ with GeCO (ours) across five tracks; higher is better. Track averages correspond to the averages over the 10 tasks in each track.
Task Track 1 Track 2 Track 3 Track 4 Track 6
$\pi_{0}$ Ours $\pi_{0}$ Ours $\pi_{0}$ Ours $\pi_{0}$ Ours $\pi_{0}$ Ours
add_condiment 0.66 0.80 0.14 0.08 0.34 0.16 0.26 0.10 0.56 0.76
insert_flower 0.18 0.30 0.04 0.00 0.22 0.28 0.02 0.00 0.10 0.02
select_book 0.694 0.84 0.064 0.08 0.417 0.51 0.311 0.23 0.714 0.79
select_chemistry_tube 0.52 0.76 0.12 0.14 0.70 0.94 0.06 0.02 0.28 0.71
select_drink 0.52 0.78 0.224 0.30 0.08 0.10 0.10 0.04 0.44 0.66
select_fruit 0.38 0.56 0.46 0.52 0.083 0.14 0.06 0.18 0.30 0.32
select_mahjong 0.25 0.58 0.02 0.04 0.125 0.12 0.12 0.06 0.02 0.25
select_painting 0.46 0.22 0.26 0.12 0.50 0.60 0.56 0.76 0.30 0.18
select_poker 0.54 0.70 0.26 0.54 0.06 0.22 0.12 0.52 0.28 0.52
select_toy 0.50 0.52 0.36 0.34 0.38 0.32 0.12 0.10 0.18 0.28
Avg_SR 0.47 0.61 0.212 0.22 0.291 0.34 0.173 0.20 0.322 0.45
Appendix C VLA Training Details.
We fine-tune $\pi_{0}$ / $\pi_{0.5}$ VLAs using the OpenPI training stack with LeRobot-format datasets. Across all benchmarks, we train for 30k gradient steps . Unless otherwise specified, weights are initialized from the corresponding OpenPI base checkpoint.
RoboTwin 2.0.
For RoboTwin 2.0, we fine-tune a separate $\pi_{0}$ model per task. Observations include three RGB views (cam_high, cam_left_wrist, cam_right_wrist) and robot proprioception (observation.state). We train for 30k steps with batch size 32 for each task.
LIBERO.
For $\pi_{0}$ on LIBERO, we fine-tune on physical-intelligence/libero datasets. We initialize from gs://openpi-assets/checkpoints/pi0_base/params and train for 30k steps with batch size 32.
For $\pi_{0.5}$ on LIBERO, we fine-tune for 30k steps with batch size 256. We use AdamW with gradient clipping (clip_gradient_norm=1.0) and an EMA of model weights (ema_decay=0.999). The learning rate follows a cosine schedule with warmup (warmup_steps=10,000, peak_lr=5e-5).
VLABench.
For the VLABench fine-tuning experiment, we choose to fine-tune $\pi_{0}$ on huggingface datasets vlabench/vlabench_primitive_ft_lerobot. We trained for 30k steps with batch size 256.
Appendix D OOD Detection Settings
ID/OOD split.
We study a controlled suite-level distribution shift on LIBERO by treating LIBERO-Goal as in-distribution (ID) and LIBERO-Spatial as out-of-distribution (OOD). We train a single policy on the 10 tasks of LIBERO-Goal and evaluate OOD detection on the 10 tasks of LIBERO-Spatial.
Episode and planning-call sampling.
We sample 500 ID episodes from LIBERO-Goal and 500 OOD episodes from LIBERO-Spatial. Each episode consists of multiple planning calls, where the policy generates an action chunk and executes a fixed number of actions before replanning. To avoid overweighting longer episodes, we compute planning-level statistics using a fixed number of planning calls per episode: we extract $J=\texttt{20}$ planning calls from each episode using random $J$ calls and treat each extracted planning call as one evaluation instance.
GeCO score (optimization-dynamics signal).
For each planning call at state $s^{(j)}$, GeCO performs iterative optimization for at most $K_{\max}=10$ steps and produces per-step norms $r_{k}=\|f_{\theta}(a_{k},s^{(j)})\|_{2}$. We use the final-step norm as the raw anomaly score for that planning call:
$\text{score}(\text{plan }j)=\|f_{\theta}(a^{(j)}_{K},s^{(j)})\|_{2},$
where $K\leq K_{\max}$ is the last executed refinement step (early termination uses the stopping step; otherwise $K=K_{\max}$).
Temporal filters and triggering rule.
We evaluate two online filters over $\{r_{k}\}_{k=1}^{K}$ within each planning call: (i) Moving Average with window $w=5$, triggering when $\tilde{r}_{k}>\tau_{\mathrm{ma}}$; (ii) Leaky Bucket defined by $b_{k}\leftarrow\max\{0,(1-\lambda)b_{k-1}+(r_{k}-\tau)_{+}\},$ triggering when $b_{k}\geq B$. We set $(\tau_{\mathrm{ma}},\tau,\lambda,B)=\texttt{(0.6, 0.5, 0, 0.7)}$ and keep them fixed across all experiments.
Baseline score (training-loss proxy).
As a baseline OOD signal, we follow the Diff-DAgger-style idea (Lee et al., 2025) of using the policy’s training objective as an uncertainty indicator. For each planning call, let $x_{0}$ denote the generated action chunk and let $c$ denote the conditioning input. We draw a single Gaussian noise sample $\epsilon\sim\mathcal{N}(0,I)$ and construct a noised input at timestep $t$ via linear interpolation
$x_{t}\;=\;(1-t)\,x_{0}+t\,\epsilon,$
where $T$ is the total number of diffusion steps used by the training objective. We then evaluate the flow-matching regression loss
$\mathcal{L}_{\mathrm{FM}}(x_{0},c)=\mathbb{E}\left\|f_{\theta}(x_{t},t,c)-(\epsilon-x_{0})\right\|_{2}^{2}$
AUROC computation.
We compute planning-level AUROC by sweeping a threshold over the scalar score to separate OOD planning calls (positive) from ID planning calls (negative), using the $500\times J$ ID instances and $500\times J$ OOD instances.
Operating point (TNR/TPR).
We report TNR/TPR at a single operating point selected to achieve TPR $\approx 90\%$ on OOD instances for our detector when possible. Specifically, we choose the threshold $\theta$ such that $\mathrm{TPR}(\theta)$ is closest to 0.9 and report the corresponding TNR.
Early reporting and Time Saved.
The detector runs online during an episode. Once any planning call triggers an OOD flag, we immediately terminate the episode. We define the reporting time $t_{\mathrm{report}}$ as the environment time step when the first trigger occurs, and compute
$\text{Time Saved}=1-\frac{t_{\mathrm{report}}}{T_{\mathrm{total}}},$
where $T_{\mathrm{total}}=300$ is the episode horizon under the standard evaluation protocol. We report Time Saved averaged over OOD episodes only.
Appendix E Real World Experiment Settings
Hardware Configuration.
We conduct our real-world experiments using the Galaxea R1 Lite mobile manipulator. The robot is equipped with two dexterous arms and 3 RGB cameras. For our policy input, we utilize three camera views: a head camera (cam_head), and two wrist-mounted cameras (cam_left_wrist, cam_right_wrist). We also include the robot’s proprioceptive state (observation.state), which contains joint positions and gripper states. The control frequency of the physical system is set to 15 Hz, and the execution action chunk size is set to $T_{a}=16$.
Data Collection and Task Description.
We select two contact-rich, high-precision manipulation tasks: Nut Assembly and Chemistry Tube Arrangement. The task setup is shown in Fig. 5.
Figure 5: Task setups for the real-world robotic deployment, showing the configurations for both the Nut Assembly and the Chemistry Tube Arrangement tasks.
  • Nut Assembly: The robot shall grasp a plastic nut, accurately align it with a threaded bolt, and perform the insertion. A trial is considered successful only when both two nuts are fully threaded onto their respective bolts.
  • Chemistry Tube Arrangement: The robot shall grasp chemistry tubes from a randomized initial pose and insert them into designated tight-tolerance slots on a rack with random position and rotation. A successful trial requires the tube to be fully seated in the correct slot without dropping or tipping over.
Training Details.
Our baseline policy is Galaxea G0 Plus, which employs a standard flow-matching-based continuous action head. To train our G0 Plus with GeCO policy, we initialize the model with the pre-trained G0 Plus weights. Following the GeCO formulation, we fix the time conditioning to $t=0$ during fine-tuning to learn a stationary velocity field. We fine-tune the model on the expert demonstrations for 4 full epochs with a batch size of 32. We use the AdamW optimizer with a peak learning rate of 2.5e-5 and a cosine learning rate schedule with warmup_steps=200.
GeCO Inference Setup.
During real-world deployment, the GeCO policy synthesizes action chunks via iterative optimization. We set the maximum number of optimization steps to $K_{\max}=10$. We use a step-size schedule $\{\eta_{k}\}$ initialized at 0.25 and decaying to 0.02 for deeper refinement. The adaptive early stopping mechanism is triggered when the gradient norm falls below a threshold $\tau_{\mathrm{opt}}$:
$\|f_{\theta}(a_{k},s_{t})\|_{2}<\tau_{\mathrm{opt}},$
where $\tau_{\mathrm{opt}}$ is empirically set to 1.5. If the threshold is not met, the optimization continues until $K_{\max}$ is reached, ensuring complex states (e.g., insertion alignment) receive maximum computational budget while simple transits exit early.
Evaluation Protocol.
We evaluate both the baseline (G0 Plus) and our method (G0 Plus + GeCO) over 10 independent physical trials for each task. The initial positions of the objects (nuts, bolts, tubes, racks) are randomized within the workspace to test robustness to spatial variations. We report the success rate and record the average Number of Function Evaluations (NFE) per planning call. For the baseline, inference is bound to a fixed schedule of NFE =10=10. For GeCO, NFE varies dynamically based on the early stopping criteria.
Appendix F Detailed Analysis of Real-World Adaptive Computation
To provide a deeper understanding of how GeCO dynamically modulates its computation over time, we present a detailed step-by-step NFE analysis of the physical rollouts in Figure 6.
Figure 6: Dynamic NFE allocation during real-world execution. We visualize the real-time Number of Function Evaluations (NFE) employed by GeCO across a full episode for both the (a) Nut Assembly and (b) Tube Arrangement tasks. Instead of relying on a fixed computation schedule, GeCO automatically allocates maximum compute (NFE=10) during critical bottleneck phases—such as precise target alignment, picking, and insertion. Conversely, the policy adaptively exits early (with NFE dropping as low as 3) during simpler free-space transit phases. This dynamic temporal modulation directly translates to the reduced average NFE (6.45 and 5.11, respectively) and improved real-time inference efficiency on physical hardware without sacrificing task success rates.
As shown in Figure 6, the high-compute peaks strictly align with the most complex physical interactions (e.g., the precise insertion phase at the end of both tasks). The ability to seamlessly transition into a low-compute regime during simpler semantic stages highlights the efficiency of our Time-Unconditional Flow Matching formulation in real-world scenarios.